OpenStudy (anonymous):

um i need sumbody smart kay ? lol bc im not ... :D what is 5 1/6 subtract 1 3/8 (you dont have to simplfy it ill do that)

OpenStudy (anonymous):

thank you if you know how to do it i got a bad bad grade im math (78) and im grounded and if i dont get my hw right umm belt?

OpenStudy (anonymous):

5 1/6 - 1 3/8 = 5.791666667

OpenStudy (anonymous):

um how do you right that as a fraction lemon ninja????????????????

OpenStudy (anonymous):

and jk bout the belt loli dont get that.. :D

OpenStudy (anonymous):

about 5 8/10 I believe

OpenStudy (anonymous):

thank you and it simplfied is 5 4/5 i think

OpenStudy (anonymous):

katieberry - you need to convert mixed fractions given (5 1/6 and 1 3/8) into improper fractions. Then subtract. 5 1/6 = 5 + 1/6 = (5*6+1)/6 = 31/6 1 3/8 = 1 + 3/8 = (1*8+3)/8 = 11/8 Now, you subtract 31/6 and 11/8.

OpenStudy (anonymous):

thank you!!! IT BETTER BE RIGHT LOL